SXT CHIEF OBSERVER CHECK LIST

SXT CO FOR WEEK

WEEKLY TASKS

Insert compressed and low8 diffuser images in a table for at least two visible orbits. Include 1/4 FL1N and 1/4 FC1N DPE=2 images (in FFI2 3-3, 3-4).

DAILY TASKS

These are tasks that should be checked on a daily basis or “as needed” basis.

M T W T F S

View http://isass1.solar.isas.ac.jp/sxt co/quick/quick.html.

Check SOHO/TRACE plans. Check E-Mail: GBO and ground-based observers. Check all the new FFI and PFI data with YODAT or RD SDA. Confirm PFI starting exposures and FFI short exposures. Run first light,/sum to check solar situation. Prepare SXT table and send Fax: ~sxt co/ksc fax Run ksc commands from IDL to archive the table Check that KSC data are being reformatted OK. Check the ARS patrol image location (if the table changed). Make sure there is paper in the LPS17 printer. Make necessary changes to ~sxt co/.mailrc file. Need to send an alert? See alerts.html.

SPECIFIC TASKS

These tasks are in addition to the “Daily Tasks.” The specific day might vary by one or two days for some tasks (e.g., orbital elements sometimes arrive a day late).

Monday

Attend Operations Meeting E-mail (alias “plan”) the WOM report and this week’s table plan (~sxt co/wom rep) E-mail (alias “status”) the SXT [bi-]weekly status message. (~sxt co/sxt status rep)

Is the SXTweekly.html updated? (/home/www/sxt co/SXTweekly.html)

Make html version of WOM report(plan2html.pro) and move to the nugget directory. Update the campaign Web files (Use sxt co account)

Tuesday

DKCAL: Run the dark calibration table

Wednesday

Check that Orbital Elements have arrived ($SSWDB/ydb/orbit sol) Is a pointing update needed (see quick.html)? Run new normal point

Thursday

See that the Visible diagram arrives from KSC (by FAX)

Run copy visible (as sxt co) and pr visible

E-mail draft plan for next week’s table load to yohkoh@flare1

Friday

Write a new nugget (/home/www/sxt co/SXTweekly.html and index.html)
Update NASA HQ (http://ossim.hq.nasa.gov/OSSIM/home.htm; yohs; update1)

Saturday

Does Monday’s table need to be sent? Prepare report for next week’s WOM (directory= wom rep)

Print checksheet for next week’s CO (~sxt co/documents/co check list.ps)

Terminator Exposures from file: term expose.tex

                     TERMINATOR EXPOSURES
                     Filter   Exposure    w/NuDen
                     ----------------------------
                     Al01        21        17
                     Mg3         23        17
                     Be          23        18
                     AlMg        25        18
                     Al12        25        18

SXT Effective Exposures DPE Exposure DPE Exposure DPE Exposure DPE Exposure 0  0.0773 ms ∗ 10  28.0 ms  20  948. ms  30  30.2 sec 1  0.232 ms ∗ 11  38.0 ms  21  1.34 sec  31  42.7 sec 2  0.960 ms  12  58.0 ms  22  1.89 sec  32  60.4 sec 3  1.37 ms ∗ 13  78.0 ms  23  2.67 sec  33  85.4 sec 4  2.88 ms  14  118. ms  24  3.78 sec  34  121. sec 5  3.06 ms ∗ 15  168. ms  25  5.34 sec  35  171. sec 6  4.67 ms ∗ 16  238. ms  26  7.55 sec  36  242. sec 7  6.28 ms ∗ 17  338. ms  27  10.7 sec 8  9.50 ms ∗ 18  468. ms  28  15.1 sec 9  17.0 ms  19  668. ms  29  21.4 sec

∗ The 8.05% mask is applied.

SXT Filter Positions Pos. Rear wheel Front wheel 1  Open  Open 2  Al.1  NaBan 3  AlMg  UV Flood 4  Be119  Diffuser 5  Al12  WdBan 6  Mg3  NuDen